Circular No. 8877

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 T1 (McNAUGHT) R. H. McNaught reports his discovery of a comet (discovery observa-

tion tabulated below) on CCD images obtained with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring survey; the object showed a 2′.0 coma, slightly extended in p.a. 80^{o}, with moderate to strong condensation. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, numerous other astrometric CCD observers have commented on the object’s cometary appearance, including L. Donato and V. Gonano, Remanzacco, Italy (0.45-m f/4.4 reflector, Oct. 9.72 UT; thirty co-added 10-s exposures obtained in twilight show a coma nearly 30′′ in diameter with a sharp central condensation and a fan tail almost 40′′ long toward p.a. 45^{o}); E. Pettarin, G. Lombardi, and L. Bittesini, Farra d’Isonzo, Italy (0.40-m f/4.5 reflector, Oct. 9.73; 25′′ diffuse coma in a poor sky); F. Hormuth, Heidelberg-Königstuhl (0.70-m reflector, Oct. 9.73–9.74; coma diameter ≈ 20′′; stacked R-band images totalling 300 s show the coma slightly elongated toward the east); M. Tichý, M. Honková, and J. Tichá, Kleť (0.57-m f/5.2 reflector, Oct. 9.76–9.77; coma diameter > 25′′); and P. Birtwhistle, Great Shefford, Berkshire, England (0.40-m f/6 Schmidt-Cassegrain reflector, Oct. 9.77–9.78; 20′′ concentrated coma with a 75′′ tail in p.a. 80^{o}).

2007 UT  α_{2000}  δ_{2000}  Mag.
Oct. 9.42315  17^{h}18^{m}17.^{s}19  −9^{o}37′17′′.5  12.8

The available astrometry, the following very preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-T54.

T = 2007 Nov. 23.742 TT  ω = 284.^{o}051
Ω = 136.922 }2000.0
q = 0.30498 AU  i = 68.327

COMET 191P/McNAUGHT Comet P/2007 N1 = P/2000 P3 (cf. IAUC 8855, 8866) has been as-

signed the permanent number 191P (cf. MPC 60617).

P/2007 R5 = 1999 R1 = 2003 R5 (SOHO) Regarding IAUC 8872, lines 16–18, the brightness peak in 2007 actually

occurred shortly before perihelion.

2007 October 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green